prove that the coordinates of the middle point of the line joining the points (2,3) and (3,4) satisfy the equation x - y + 1=o
1
Expert's answer
2013-08-20T11:12:35-0400
Calculate coordinates of the middle point: x= (2 + 3)/2 = 2.5, y = (3 + 4)/2 = 3.5.
Middle point is (2.5, 3.5).
Then we substitute coordinates of middle point in equation x - y + 1 = 0:
2.5 - 3.5 + 1 = -1 + 1 = 0 that proves our claim.